OpenStudy (anonymous):

A certain freely falling object, released from rest, requires 1.85 s to travel the last 33.0 m before it hits the ground. (a) Find the velocity of the object when it is 33.0 m above the ground. (Indicate the direction with the sign of your answer. Let the positive direction be upward.) (b) Find the total distance the object travels during the fall. I got 8.8 m/s for Part A but can't seem to get the correct answer for Part B (which I got 7.6 m).

OpenStudy (anonymous):

From your answer to part A, you can calculate the total time the object was falling: \[V_f = v_a - g * \Delta t \] Substituting in the results: \[V_f = -8.8 -10*1.85 = -27.3 (m/\sec)\] Since the object is released from rest: \[V_f = 0 -g*t_T\] \[-27.3 = -10*t_T\] \[t_T = 2.73\] And now you can plug in the total falling time into y(t) eq. : \[y(t) = -0.5*g*t_T^2\] \[y(t_T) = -5*2.73^2 = 37.26(m)\]